OpenStudy (anonymous):

given that f(x)=x^6h(x) h(−1)=2 h′(−1)=5 f′(−1)=?

OpenStudy (anonymous):

Is it f(x)=x^(6h(x)) or f(x)=(x^6)h(x)?

OpenStudy (anonymous):

x^(6)h(x)

OpenStudy (anonymous):

Alright, then. We know that f'(x)=h'(x)(x^6)+(6x^5)h(x). At x=-1, f(-1)=((-1)^6)h(-1)=1*2 (we're given the solution for h(-1)). We want to find f'(-1)=h'(-1)((-1)^6)+(6(-1)^5)h(-1). From here, we can solve it.

OpenStudy (anonymous):

you mind if i ask how?

OpenStudy (anonymous):

Plug things in. We know h'(-1), as we're given it. We know h(-1), since we also are given that.

OpenStudy (anonymous):

So I am plugging -1 into both equations and doing what with the result? Multiplying?

OpenStudy (anonymous):

We want to find f'(-1), right? Taking the derivative of f(x) should give us f'(x). So I took the derivative. f'(x)=h'(x)(x^6)+(6x^5)h(x) Now plug in -1 for x. So we have f'(-1)=h'(-1)((-1)^6)+(6(-1)^5)h(-1). By solving the right side, which we can since we know the values for all the terms, we can find f'(-1).

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

"Do or do not. There is no try." ~ Yoda

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

For my answer, I received -175

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

I received -7. At this point it's basic calculation mistakes. You should check your math.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you help me with this one if you don't mind?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

The graphs of the function f (given in blue) and g (given in red) are plotted above. Suppose that u(x)=f(x)g(x) and v(x)=f(x)/g(x). Find each of the following: u'(1) v'(1)

OpenStudy (anonymous):

Alright, since we're primarily concerned with the functions at 0<x<3, we should just try to model at those points alone. f(x)=x, and g(x)=-x+4. These are evident by looking at the graphs. u(x)=f(x)g(x)=-(x^2)+4x, u'(x)=-2x+4, u'(1)=2 v(x)=f(x)/g(x)=x/(-x+4), v'(x)=1/(-x+4)+x/(-x+4)^2, v'(1)=(1/3)+(1/9)=4/9 I think my math is right, but you should follow these steps too to make sure.

OpenStudy (anonymous):

last one x 11 1 2 −3 f(x) −1209 1 −3 37 g(x) 2432 2 11 −74 f′(x) −341 −1 −8 −33 g′(x) 683 3 17 67 (fg)′(2) f(2)/(g(2)+5) (fg)(1) (f−g)′(2)

OpenStudy (anonymous):

Does it say (fg)'(2) for the first one? I don't think that makes sense. Typo?

OpenStudy (anonymous):

You sure it isn't (f(g'(x)) or something?

OpenStudy (anonymous):

its (fg)'2

OpenStudy (anonymous):

This is a markup I'm unfamiliar with. Sorry, can't help you here.

OpenStudy (anonymous):

what about the rest